Find the area of the curve y2 (2a – x) = x3
between the area and its asymptotes.
1
Expert's answer
2021-07-20T08:12:00-0400

The asymptote(s) of the curve parallel to "y" -axis is given by "2a-x=0."

Then the vertical asymptote is "x=2a."


"y=\sqrt{\dfrac{x^3}{2a-x}}"

Let "x=2a\sin^2\theta."



"dx=4a\sin \theta \cos \theta d\theta"

"Area=A=\displaystyle\int_{0}^{2a}\sqrt{\dfrac{x^3}{2a-x}}dx"

"=\displaystyle\int_{0}^{\pi/2}\sqrt{\dfrac{8a^3\sin^6\theta}{2a-2a\sin^2\theta}}4a\sin \theta \cos \theta d\theta"

"=8a^2\displaystyle\int_{0}^{\pi/2}\dfrac{\sin^3\theta\sin \theta \cos \theta}{\cos \theta}d\theta"

"=8a^2\displaystyle\int_{0}^{\pi/2}\sin^4\theta d\theta=2a^2\displaystyle\int_{0}^{\pi/2}(1-\cos(2\theta))^2 d\theta"

"=2a^2\displaystyle\int_{0}^{\pi/2}(1-2\cos(2\theta)+\cos^2(2\theta)) d\theta"

"=a^2\displaystyle\int_{0}^{\pi/2}(2-4\cos(2\theta)+1+\cos(4\theta)) d\theta"

"=a^2\bigg[3\theta-2\sin(2\theta)+\dfrac{1}{4}\sin4\theta()\bigg]\begin{matrix}
 \pi/2 \\
 0
\end{matrix}"

"=\dfrac{3\pi a^2}{2} (units^2)"

"Area=\dfrac{3\pi ^2}{2}" square units.